# Accelerated Ostwald ripening by chemical activity

### v1

@@ Acknowledgements
SII.3 Gas phase
Since the gas is dilute, we claimed in the main text that $D_{ij\mathrm{g}}=\delta_{ij}D_{\mathrm{g}}$, where $D_{\mathrm{g}}$ is an effective single-particle diffusion constant. This is the case since both the hydrodynamic and thermodynamic couplings are weak as the sticky molecules rarely encounter each other. This is further supported by inserting the FH free energy Eq. (S28) and the mobility Eq. (S41) in Eq. (S40a),
@@
We proceed to investigate the reactions, $\partial\bar{\phi}(t)/\partial t=\bar{\upsilon}s(\bar{\phi}(t))$. Once a total amount of material is imposed (e.g., in our case, $\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\varphi}^{\infty}=\varphi^{\infty}$), chemical equilibrium is determined by the condition $s(\bar{\varphi}^{\infty})=0$. For it to be a stable equilibrium, we require that the Jacobian, $\bar{\upsilon}(\partial s/\partial\bar{\phi})^{\mathrm{T}}|_{\bar{\varphi}^{\infty}}\simeq\bar{\upsilon}\bar{k}_{\mathrm{g}}^{\mathrm{T}}$ does not have any positive eigenvalues; otherwise, if the volume fractions deviate from $\bar{\varphi}^{\infty}$ (which is indeed the case around a surface, where the volume fractions are slightly different, $\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$), the entire dilute phase would destabilize due to the mere presence of reactions. We make the reasonable assumption that the reactions in isolation are indeed stable. The Jacobian’s eigenvalues are 00 (corresponding to steady state) and $\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\upsilon}$, so we require $\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\upsilon}<0$ (corresponding to a decaying mode). Since $\bar{\bar{D}}_{\mathrm{g}}$ is to leading order diagonal, we confirm that $\ell^{2}_{\mathrm{g}}\simeq-D_{\mathrm{g}}\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\upsilon}>0$ (as $D_{\mathrm{g}}>0$), so the choice of sign in Eq. (S44) is justified.
SII.3.1 Solution of the linearized equations
@@
SII.3.2 Simpler main-text linearization
The results of the previous section stem from a general linearization procedure, and they will be used in Sec. SII.5 below. Owing to the gas phase being dilute, the results of the main text were based on a slightly simpler procedure as follows: The volume fraction fields $\bar{\phi}_{\mathrm{g}}(r)$ vary between (the small) $\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$ and (the small) $\bar{\varphi}^{\infty}$, so they are small for all $r$. Thus, we expand the reaction rates around zero volume fraction (as opposed to around $\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$ as above), $s(\bar{\phi})\simeq s(\bar{0})+(\partial s/\partial\bar{\phi})^{\mathrm{T}}|_{\bar{0}}\bar{\phi}$. There are no reactions in the absence of material, so $s(\bar{0})=0$. Next, we assume that there is a first-order reaction step in the interconversion mechanism, such that neither component of $(\partial s/\partial\bar{\phi})^{\mathrm{T}}|_{\bar{0}}$ is zero. Hence, we neglect higher-order terms in the above expansion. For first-order reactions to take place, the stoichiometry must be one to one, $\bar{\upsilon}=(-1,1)^{\mathrm{T}}$. Furthermore, we denote as in the main text, $(\partial s/\partial\bar{\phi})|_{\bar{0}}=(k_{\mathrm{s}\to\mathrm{ns}},-k_{\mathrm{ns}\to\mathrm{s}})^{\mathrm{T}}$ (with constant $k_{\mathrm{s}\to\mathrm{ns}},k_{\mathrm{ns}\to\mathrm{s}}>0$), where the sign is such that $\mathrm{s}$ is lost when creating $\mathrm{ns}$ and vice versa. Overall,
@@
which was used on the right-hand side of Eq. (15) in the main text. This approximation simply states that only the first-order reactions occur with a meaningful rate, as encounters between molecules required for higher-order reactions are too rare. This need not imply necessarily that Eq. (S50) is the full mechanism; a different limit of the full mechanism might apply in the liquid phase.
In the language of the more general linearization of Sec. SII.2, the main-text linearization amounts to choosing $\bar{k}_{\mathrm{g}}=(k_{\mathrm{s}\to\mathrm{ns}},-k_{\mathrm{ns}\to\mathrm{s}})^{\mathrm{T}}$ and $s_{\mathrm{g}}=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{L}$. Using the diagonal diffusivity $D_{ij\mathrm{g}}=\delta_{ij}D_{\mathrm{g}}$ and the stoichiometry $\bar{\upsilon}=(-1,1)^{\mathrm{T}}$, we obtain $1/\ell_{\mathrm{g}}^{2}=(|\bar{\upsilon}_{\mathrm{s}}|k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})/D_{\mathrm{g}}$ (which is Eq. (19) of the main text). Upon inserting these facts in Eqs. (S47), we find
@@
which coincide with Eqs. (17) and (18) of the main text. Explicitly,
@@
where $\bar{\varphi}^{\infty}=(k_{\mathrm{ns}\to\mathrm{s}}\varphi^{\infty}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}),k_{\mathrm{s}\to\mathrm{ns}}\varphi^{\infty}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}))^{\mathrm{T}}$. As mentioned earlier, the first two terms are the standard diffusive profile around a droplet, whereas the last terms oppose departure from the chemical equilibrium, increasing the diffusive flux out of small droplets and into large droplets. To see that explicitly, the fluxes from Eq. (S49) simplify as
@@
In the derivations to follow, we use the more general linearized result of Sec. SII.3.1. It accounts for cases where a higher-order reaction dominates, namely, if $(\partial s/\partial\bar{\phi})|_{\bar{0}}=\bar{0}$. For example, $s(\bar{\phi})=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}}^{2}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}}^{2}$ with $\bar{\upsilon}=(-1,1)$, which describes a reaction mechanism where a collision between two particles of the same form is needed to form two of the other form. In that case, one should have continued the above series to $s(\bar{\phi})\simeq\bar{\phi}^{\mathrm{T}}(\partial^{2}s/\partial\bar{\phi}\partial\bar{\phi}^{\mathrm{T}})|_{\bar{0}}\bar{\phi}$, such that $\bar{k}_{\mathrm{g}}^{\mathrm{T}}=2\bar{\phi}_{\mathrm{g}}^{\mathrm{L}\mathrm{T}}(\partial^{2}s/\partial\bar{\phi}\partial\bar{\phi}^{\mathrm{T}})|_{\bar{0}}$ and $s_{\mathrm{g}}=\bar{\phi}_{\mathrm{g}}^{\mathrm{L}\mathrm{T}}(\partial^{2}s/\partial\bar{\phi}\partial\bar{\phi}^{\mathrm{T}})|_{\bar{0}}\bar{\phi}_{\mathrm{g}}^{L}=(1/2)\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$ (as opposed to the above, effectively first-order treatment where $s_{\mathrm{g}}=\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$). In total, the general linearization of Sec. SII.2 accounts for non-first-order reaction mechanisms, which is a slightly-more general formalism than the one in the main text.
SII.4 Liquid phase
@@
where we have kept up to first-order corrections in small $e^{-\chi}\ll 1$ and $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}\ll 1$ and we omitted Laplace-pressure corrections for simplicity. This matrix incorporates both the mobility and the free-energy response to fluctuations in the volume fraction fields. Since the liquid phase is dense in sticky protein, the sticky proteins quickly ($\sim\Gamma e^{\chi}$) rearrange in response to a perturbation in either the sticky or nonsticky volume fraction. (At the same time, the nonsticky protein is inert, so $D_{\mathrm{ns}\mathrm{s}\mathrm{l}}\sim 0$ and $D_{\mathrm{ns}\mathrm{ns}\mathrm{l}}\sim 1$ as if it is still an ideal gas.) This dominant order-$e^{\chi}$ term arises from the second-order derivative of the entropy contribution to the FH free energy, as the sticky volume fraction is very close to $1$ (within $\sim e^{-\chi}$) and so approaching it further is entropically costly. Interestingly, since the sticky proteins attract each other, a sticky fluctuation is slightly more stable than a nonsticky fluctuation (cf. $D_{\mathrm{s}\mathrm{s}\mathrm{l}}-D_{\mathrm{s}\mathrm{ns}\mathrm{l}}=-\Gamma(2\chi-1)<0$ with $\chi\gg 1$) as the former is slightly stabilized by the energy term. This will play a role in determining $\ell_{\mathrm{l}}^{2}$ in Sec. SII.4.3. Generally, the liquid-phase analogue for the reaction-stability criterion ($\bar{k}_{\mathrm{l}}^{\mathrm{T}}\bar{\upsilon}<0$) does not impose anything on $\ell_{\mathrm{l}}^{2}=-\bar{k}_{\mathrm{l}}^{\mathrm{T}}\bar{\bar{D}}_{\mathrm{l}}^{-1}\bar{\upsilon}$ since $\bar{\bar{D}}_{\mathrm{l}}$ is not diagonal. We proceed to Secs. SII.4.1 and SII.4.2 as if the sign convention in Eq. (S44) is correct; this will not be the case in Sec. SII.4.3.
SII.4.1 Solution of the linearized equations
@@
In Secs. SII.4.2 and SII.4.3, we examine the validity of the linearization approximation, and find two destabilizing effects that one must avoid in order to consider the possibility of acceleration.
@@
SII.4.2 Example: Crossing the spinodal
@@
In contrast to the gas phase (Sec. SII.3.2), $s_{\mathrm{l}}\sim 1\gg s_{\mathrm{g}}\sim\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$. As a result, we see in Eq. (S62) that the deviations from the coexistence (binodal) volume fractions $\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}$ might be of order up to $k_{\mathrm{s}\to\mathrm{ns}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\sim 1$. When the droplets are small ($R\ll\ell_{\mathrm{l}}$), the second term would be of order $[k_{\mathrm{s}\to\mathrm{ns}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})](R^{2}-r^{2})/\ell_{\mathrm{l}}^{2}$ which is small ($R^{2}/\ell_{\mathrm{l}}^{2}\ll 1$), so the linearization approach is indeed justified. On the other hand, for a large droplet, the volume fractions at the core are modified by the above factor $k_{\mathrm{s}\to\mathrm{ns}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})$. Therefore, if reactions favor the formation of nonsticky species ($k_{\mathrm{s}\to\mathrm{ns}}>k_{\mathrm{ns}\to\mathrm{s}}$, which is what is needed to obtain a meaningful acceleration), the second term in Eq. (S62) originating from the reactions becomes of order $\sim 1$. This (i) invalidates the linearization approach, and (ii) even worse, suggests that if reactions dominate over the compressibility, this might drive the volume fractions beyond the spinodal, causing the whole dense-phase to destabilize. We recall that in this situation, where $s_{\mathrm{l}}\sim\Gamma/\ell_{\mathrm{l}}^{2}\gg\Gamma(\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}})/\ell_{\mathrm{l}}^{2}$, we concluded in the main text (Sec. IV A) that size control takes place [3], and hence this instability is averted. Although, no acceleration would be achieved in this case.
SII.4.3 Example: Turing instability
@@
These expressions, combined with the diffusion matrix of Eq. (S48), are inserted in Eq. (S44) yield $1/\ell_{\mathrm{l}}^{2}=e^{-\chi}[(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})/\Gamma][1-2\chi k_{\mathrm{s}\to\mathrm{ns}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})]$. This expression can be understood as follows: First, the combination of very little access for interconversion enzymes into droplets ($s\sim 1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}}\ll 1$) with the rapid relaxation of fluctuations in sticky volume fraction field ($D_{\mathrm{s}i\mathrm{l}}\sim e^{\chi}$), implies that the typical distance between reactions is large, $\ell_{\mathrm{l}}\sim e^{\chi/2}$. Second, $\ell_{\mathrm{l}}^{2}$ in its current definition (Eq. (S44)) might become negative if $\chi>(1+k_{\mathrm{ns}\to\mathrm{s}}/k_{\mathrm{s}\to\mathrm{ns}})/2$, i.e., if the nonsticky protein is not converted to the sticky form fast enough. This is a result of a Turing instability, whereby diffusion destabilizes the otherwise stable reactions ($(\partial s/\partial\bar{\phi})^{\mathrm{T}}\bar{\upsilon}<0$). Specifically, as we identified in Eq. (S54), the entropic response of the sticky volume-fraction field to a fluctuation in the nonsticky one is faster than the response to a fluctuation in the sticky one (since the latter is slightly stabilized energetically, parametrized with $\chi$). Thus, unless nonsticky proteins are converted back fast enough, the sticky proteins would separate from regions where nonsticky proteins are created. Since surface thermodynamics maintain $\phi_{\mathrm{s}\mathrm{l}}(R)=\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}$ at the interface, this instability would start at the core of the droplet every distance “$\mathrm{i}\ell_{\mathrm{l}}$”. Due to the surface tension, we might expect that this would culminate in bubbles [9].
@@
Indeed, when the droplets are larger than the Turing length, $R>\pi\ell_{\mathrm{T}}$, the instabilities start forming, and according to Eq. (S65) the volume-fraction fields found from the linear theory exhibit divergence. As in Sec. SII.4.2, this indicates that the linear theory is no longer valid, and the volume fraction fields might cross the spinodal to form “dilute pockets” inside the droplets, whose wavelength is $1/\ell_{\mathrm{T}}$.
However, for small droplets ($R\ll\ell_{\mathrm{T}}$), the patterns are yet to develop, and instead the volume fractions are roughly uniform inside the droplet, as in Sec. SII.4.1. Replacing $-\ell_{\mathrm{l}}^{2}\to\ell_{\mathrm{T}}^{2}$ in Eq. (S48), the previously-obtained flux (Eq. (S57)) becomes instead
@@
SII.4.4 Linearization-free treatment
If the reactions destabilize the liquid phase (whether due to the passing of the spinodal line or through a Turing-instability mechanism), discussing accelerated ripening becomes meaningless. Thus, we assume that neither of these effects occurs, and the free energy is able to oppose any reaction-related instabilities. Overall, we are motivated to avoid the linearization in the liquid phase.
We revert back to the quasi-static approximation (see Sec. SII.2). With spherical symmetry, our starting point is the steady-state equation for the volume-fraction field,
@@
where $J_{\mathrm{l}}(r)$ is the (unlinearized) steady-state flux within the dense phase for a fixed radius $R$, and the boundary conditions are given in Eqs. (S55). We multiply both sides by $r^{2}$ and integrate from 00 to $R$, which results in the general expression

### v2

@@ Acknowledgements
SII.4 Gas phase
Since the gas is dilute, we consider a diagonal matrix $D_{ij\mathrm{g}}=\delta_{ij}D_{\mathrm{g}}$, where $D_{\mathrm{g}}$ is the effective single-particle diffusion constant that appears in the main text. This is reasonable since both the hydrodynamic and thermodynamic couplings are weak as the sticky molecules rarely encounter each other. This is further supported by inserting the FH free energy Eq. (S29) and the mobility Eq. (S45) in Eq. (S44a),
@@
We proceed to investigate the reactions, $\partial\bar{\phi}(t)/\partial t=\bar{\upsilon}s(\bar{\phi}(t))$. Once a total amount of material is imposed (e.g., in our case, $\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\varphi}^{\infty}=\varphi^{\infty}$), chemical equilibrium is determined by the condition $s(\bar{\varphi}^{\infty})=0$. For it to be a stable equilibrium, we require that the Jacobian, $\bar{\upsilon}(\partial s/\partial\bar{\phi})^{\mathrm{T}}|_{\bar{\varphi}^{\infty}}\simeq\bar{\upsilon}\bar{k}_{\mathrm{g}}^{\mathrm{T}}$ does not have any positive eigenvalues; otherwise, if the volume fractions deviate from $\bar{\varphi}^{\infty}$ (which is indeed the case around a surface, where the volume fractions are slightly different, $\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$), the entire dilute phase would destabilize due to the mere presence of reactions. We make the reasonable assumption that the reactions in isolation are indeed stable. The Jacobian’s eigenvalues are 0 (corresponding to steady state) and $\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\upsilon}$, so we require $\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\upsilon}<0$ (corresponding to a decaying mode). Since $\bar{\bar{D}}_{\mathrm{g}}$ is to leading order diagonal, we confirm that $\ell^{2}_{\mathrm{g}}\simeq-D_{\mathrm{g}}\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\upsilon}>0$ (as $D_{\mathrm{g}}>0$), so the choice of sign in Eq. (S48) is justified.
SII.4.1 Solution of the linearized equations
@@
SII.4.2 Simpler main-text linearization
The results of the previous section stem from a general linearization procedure, and they will be used in Sec. SII.6 below. Owing to the gas phase being dilute, the results of the main text were based on a slightly simpler procedure as follows: The volume fraction fields $\bar{\phi}_{\mathrm{g}}(r)$ vary between (the small) $\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$ and (the small) $\bar{\varphi}^{\infty}$, so they are small for all $r$. Thus, we expand the reaction rates around zero volume fraction (as opposed to around $\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$ as above), $s(\bar{\phi})\simeq s(\bar{0})+(\partial s/\partial\bar{\phi})^{\mathrm{T}}|_{\bar{0}}\bar{\phi}$. There are no reactions in the absence of material, so $s(\bar{0})=0$. Next, we assume that there is a first-order reaction step in the interconversion mechanism, such that neither component of $(\partial s/\partial\bar{\phi})^{\mathrm{T}}|_{\bar{0}}$ is zero. Hence, we neglect higher-order terms in the above expansion. For first-order reactions to take place, the stoichiometry must be one to one, $\bar{\upsilon}=(-1,1)^{\mathrm{T}}$. Furthermore, we denote as in the main text, $(\partial s/\partial\bar{\phi})|_{\bar{0}}=(k_{\mathrm{s}\to\mathrm{ns}},-k_{\mathrm{ns}\to\mathrm{s}})^{\mathrm{T}}$ (with constant $k_{\mathrm{s}\to\mathrm{ns}},k_{\mathrm{ns}\to\mathrm{s}}>0$), where the sign is such that $\mathrm{s}$ is lost when creating $\mathrm{ns}$ and vice versa. Indeed, $k_{i\to j}$ are uniquely determined from the full reaction mechanism $s(\bar{\phi})$. Overall,
@@
which was used for the gas-phase reactions in the main text. This approximation simply states that only the first-order reactions occur with a meaningful rate, as encounters between molecules required for higher-order reactions are too rare. This need not imply necessarily that Eq. (S54) is the full mechanism; a different limit of the full mechanism might apply in the liquid phase.
In the language of the more general linearization of Sec. SII.3, the main-text linearization amounts to choosing $\bar{k}_{\mathrm{g}}=(k_{\mathrm{s}\to\mathrm{ns}},-k_{\mathrm{ns}\to\mathrm{s}})^{\mathrm{T}}$ and $s_{\mathrm{g}}=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{L}$. Using the diagonal diffusivity $D_{ij\mathrm{g}}=\delta_{ij}D_{\mathrm{g}}$ and the stoichiometry $\bar{\upsilon}=(-1,1)^{\mathrm{T}}$, we obtain $1/\ell_{\mathrm{g}}^{2}=(|\bar{\upsilon}_{\mathrm{s}}|k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})/D_{\mathrm{g}}$. Upon inserting these facts in Eqs. (S51), we find
@@
Explicitly,
@@
where $\bar{\varphi}^{\infty}=(k_{\mathrm{ns}\to\mathrm{s}}\varphi^{\infty}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}),k_{\mathrm{s}\to\mathrm{ns}}\varphi^{\infty}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}))^{\mathrm{T}}$. These are the profiles sketched in Fig. 2 of the main text. As mentioned earlier, the first two terms are the standard diffusive profile around a droplet, whereas the last terms oppose departure from the chemical equilibrium, increasing the diffusive flux out of small droplets and into large droplets. To see that explicitly, the fluxes from Eq. (S53) simplify as
@@
In the derivations to follow, we use the more general linearized result of Sec. SII.4.1. It accounts for cases where a higher-order reaction dominates, namely, if $(\partial s/\partial\bar{\phi})|_{\bar{0}}=\bar{0}$. For example, $s(\bar{\phi})=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}}^{2}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}}^{2}$ with $\bar{\upsilon}=(-1,1)$, which describes a reaction mechanism where a collision between two particles of the same form is needed to form two of the other form. In that case, one should have continued the above series to $s(\bar{\phi})\simeq\bar{\phi}^{\mathrm{T}}(\partial^{2}s/\partial\bar{\phi}\partial\bar{\phi}^{\mathrm{T}})|_{\bar{0}}\bar{\phi}$, such that $\bar{k}_{\mathrm{g}}^{\mathrm{T}}=2\bar{\phi}_{\mathrm{g}}^{\mathrm{L}\mathrm{T}}(\partial^{2}s/\partial\bar{\phi}\partial\bar{\phi}^{\mathrm{T}})|_{\bar{0}}$ and $s_{\mathrm{g}}=\bar{\phi}_{\mathrm{g}}^{\mathrm{L}\mathrm{T}}(\partial^{2}s/\partial\bar{\phi}\partial\bar{\phi}^{\mathrm{T}})|_{\bar{0}}\bar{\phi}_{\mathrm{g}}^{L}=(1/2)\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$ (as opposed to the above, effectively first-order treatment where $s_{\mathrm{g}}=\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$). In total, the general linearization of Sec. SII.3 accounts for non-first-order reaction mechanisms, which is a more general gas-phase formalism than the one in the main text.
SII.5 Liquid phase
@@
where we have kept up to first-order corrections in small $e^{-\chi}\ll 1$ and $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}\ll 1$ and we omitted Laplace-pressure corrections for simplicity. This matrix incorporates both the mobility and the free-energy response to fluctuations in the volume fraction fields. Since the liquid phase is dense in sticky protein, the sticky proteins quickly ($\sim\Gamma e^{\chi}$) rearrange in response to a perturbation in either the sticky or nonsticky volume fraction. (At the same time, the nonsticky protein is inert, so $D_{\mathrm{ns}\mathrm{s}\mathrm{l}}\sim 0$ and $D_{\mathrm{ns}\mathrm{ns}\mathrm{l}}\sim 1$ as if it is still an ideal gas.) This dominant order-$e^{\chi}$ term arises from the second-order derivative of the entropy contribution to the FH free energy, as the sticky volume fraction is very close to $1$ (within $\sim e^{-\chi}$) and so approaching it further is entropically costly. Interestingly, since the sticky proteins attract each other, a sticky fluctuation is slightly more stable than a nonsticky fluctuation (cf. $D_{\mathrm{s}\mathrm{s}\mathrm{l}}-D_{\mathrm{s}\mathrm{ns}\mathrm{l}}=-\Gamma(2\chi-1)<0$ with $\chi\gg 1$) as the former is slightly stabilized by the energy term. This will play a role in determining $\ell_{\mathrm{l}}^{2}$ in Sec. SII.5.3. Generally, the liquid-phase analogue for the reaction-stability criterion ($\bar{k}_{\mathrm{l}}^{\mathrm{T}}\bar{\upsilon}<0$) does not impose anything on $\ell_{\mathrm{l}}^{2}=-\bar{k}_{\mathrm{l}}^{\mathrm{T}}\bar{\bar{D}}_{\mathrm{l}}^{-1}\bar{\upsilon}$ since $\bar{\bar{D}}_{\mathrm{l}}$ is not diagonal. We proceed to Secs. SII.5.1 and SII.5.2 as if the sign convention in Eq. (S48) is correct; this will not be the case in Sec. SII.5.3.
SII.5.1 Solution of the linearized equations
@@
In Secs. SII.5.2 and SII.5.3, we examine the validity of the linearization approximation, and find two destabilizing effects that one must avoid in order to consider the possibility of acceleration.
@@
SII.5.2 Example: Crossing the spinodal
@@
In contrast to the gas phase (Sec. SII.4.2), $s_{\mathrm{l}}\sim 1\gg s_{\mathrm{g}}\sim\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$. As a result, we see in Eq. (S66) that the deviations from the coexistence (binodal) volume fractions $\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}$ might be of order up to $k_{\mathrm{s}\to\mathrm{ns}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\sim 1$. When the droplets are small ($R\ll\ell_{\mathrm{l}}$), the second term would be of order $k_{\mathrm{s}\to\mathrm{ns}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})](R^{2}-r^{2})/\ell_{\mathrm{l}}^{2}$ which is small ($R^{2}/\ell_{\mathrm{l}}^{2}\ll 1$), so the linearization approach is indeed justified. On the other hand, for a large droplet, the volume fractions at the core are modified by the above factor $k_{\mathrm{s}\to\mathrm{ns}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})$. Therefore, if reactions favor the formation of nonsticky species ($k_{\mathrm{s}\to\mathrm{ns}}>k_{\mathrm{ns}\to\mathrm{s}}$, which is what is needed to obtain a meaningful acceleration), the second term in Eq. (S66) originating from the reactions becomes of order $\sim 1$. This (i) invalidates the linearization approach, and (ii) even worse, suggests that if reactions dominate over the compressibility, this might drive the volume fractions beyond the spinodal, causing the whole dense-phase to destabilize. We recall that in this situation, where $s_{\mathrm{l}}\sim\Gamma/\ell_{\mathrm{l}}^{2}\gg\Gamma(\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}})/\ell_{\mathrm{l}}^{2}$ (where $\Gamma=D_{\mathrm{g}}$), we concluded in the main text that size control takes place [3], and hence this instability is averted. Although, no acceleration would be achieved in this case.
SII.5.3 Example: Turing instability
@@
These expressions, combined with the diffusion matrix of Eq. (S52), are inserted in Eq. (S48) yield $1/\ell_{\mathrm{l}}^{2}=e^{-\chi}[(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})/\Gamma][1-2\chi k_{\mathrm{s}\to\mathrm{ns}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})]$. This expression can be understood as follows: First, the combination of very little access for interconversion enzymes into droplets ($s\sim 1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}}\ll 1$) with the rapid relaxation of fluctuations in sticky volume fraction field ($D_{\mathrm{s}i\mathrm{l}}\sim e^{\chi}$), implies that the typical distance between reactions is large, $\ell_{\mathrm{l}}\sim e^{\chi/2}$. Second, $\ell_{\mathrm{l}}^{2}$ in its current definition (Eq. (S48)) might become negative if $\chi>(1+k_{\mathrm{ns}\to\mathrm{s}}/k_{\mathrm{s}\to\mathrm{ns}})/2$, i.e., if the nonsticky protein is not converted to the sticky form fast enough. This is a result of a Turing instability, whereby diffusion destabilizes the otherwise stable reactions ($(\partial s/\partial\bar{\phi})^{\mathrm{T}}\bar{\upsilon}<0$). Specifically, as we identified in Eq. (S58), the entropic response of the sticky volume-fraction field to a fluctuation in the nonsticky one is faster than the response to a fluctuation in the sticky one (since the latter is slightly stabilized energetically, parametrized with $\chi$). Thus, unless nonsticky proteins are converted back fast enough, the sticky proteins would separate from regions where nonsticky proteins are created. Since surface thermodynamics maintain $\phi_{\mathrm{s}\mathrm{l}}(R)=\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}$ at the interface, this instability would start at the core of the droplet every distance $\mathrm{i}\ell_{\mathrm{l}}$”. Due to the surface tension, we might expect that this would culminate in bubbles [9].
@@
Indeed, when the droplets are larger than the Turing length, $R>\pi\ell_{\mathrm{T}}$, the instabilities start forming, and according to Eq. (S69) the volume-fraction fields found from the linear theory exhibit divergence. As in Sec. SII.5.2, this indicates that the linear theory is no longer valid, and the volume fraction fields might cross the spinodal to form “dilute pockets” inside the droplets, whose wavelength is $1/\ell_{\mathrm{T}}$.
However, for small droplets ($R\ll\ell_{\mathrm{T}}$), the patterns are yet to develop, and instead the volume fractions are roughly uniform inside the droplet, as in Sec. SII.5.1. Replacing $\ell_{\mathrm{l}}^{2}\to\ell_{\mathrm{T}}^{2}$ in Eq. (S52), the previously-obtained flux (Eq. (S61)) becomes instead
@@
SII.5.4 Linearization-free treatment
If the reactions destabilize the liquid phase (whether due to the passing of the spinodal line or through a Turing-instability mechanism), discussing accelerated ripening becomes meaningless. Thus, we assume that neither of these effects occurs, and the free energy opposes any reaction-related instabilities. With this assumption, we proceed to implicitly find the dense-phase fluxes without linearizing the volume-fraction field equations.
We revert back to the quasi-static approximation (to before Sec. SII.3). With spherical symmetry, our starting point is the steady-state equation for the volume-fraction field,
@@
where $J_{\mathrm{l}}(r)$ is the (unlinearized) steady-state flux within the dense phase for a fixed radius $R$, and the boundary conditions are given in Eqs. (S59). We multiply both sides by $r^{2}$ and integrate from 0 to $R$, which results in the general expression